OpenStudy (anonymous):

The ordered pairs (2, –17) and (5, –35) are solutions to which of the following equations? (1 point) • y = –6x + 5 • y = 6x – 5 • y = 6x + 5 • y = –6x – 5

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (-35 - (-17) / (5 - 2) slope(m) = (-35 + 17) / 3 slope(m) = - 18/3 slope(m) = -6 knowing that the slope is -6...that eliminates the middle 2 equations now we use y = mx + b slope(m) = -6 use either of your points (2,-17)...x = 2 and y = -17 now we sub and find b -17 = -6(2) + b -17 = -12 + b -17 + 12 = b - 5 = b equation is : y = -6x - 5

OpenStudy (texaschic101):

any questions at all ?

OpenStudy (anonymous):

Thanks!!